OpenStudy (anonymous):

Identify the correct slope and y-intercept of the equation 5x - 3y = -6

OpenStudy (icalibear):

if the formula for slope is y - mx + b, what is the slope for this equation

OpenStudy (anonymous):

slope = -5; y-intercept = (0, -6)

OpenStudy (icalibear):

y = mx + b* I meant but good :)

OpenStudy (anonymous):

slope =5/3 ; y-intercept = (0, -6)

OpenStudy (icalibear):

How did you get 5/3?

OpenStudy (anonymous):

3/5

OpenStudy (icalibear):

ok nevermind sorry lol yes that's right just had to check it.

OpenStudy (icalibear):

if you ever come across this again you can use this as a reminder http://www.algebrahouse.com/7/post/2011/11/find-slope-and-y-intercept.html

OpenStudy (anonymous):

so its slope =5/3 ; y-intercept = (0, -6)

OpenStudy (anonymous):

@icalibear

OpenStudy (icalibear):

Yeah :)